x= P(x) =
0 0.68
1 0.2
2 0.06
3 0.04
4 0.02
Find P(X>1). Ans: 0.12. I don't understand how to find the P. Shouldn't it be zero?

Found 2 solutions by Boreal, math_tutor2020:
Answer by Boreal(15235)

It is not the probability is greater than 1, but rather that P(x>1)=1-P(x=0)-P(x=1)=1-0.68-0.2,and that is 0.12.

Answer by math_tutor2020(3835)


Given table:
XP(X)
00.68
10.2
20.06
30.04
40.02

P(X>1) is asking you to add up all of the probabilities in the second column that correspond to X values larger than 1.

So we'll add up these probabilities: {0.06, 0.04, 0.02} since they correspond to X = 2 through X = 4.

0.06+0.04+0.02 = 0.12

That's why P(X>1) = 0.12
There's a 12% chance of X being larger than 1.

As the tutor Boreal pointed out, you can take a (slight) shortcut in adding the probabilities for X = 0 and X = 1, then subtracting that sum from 1. This is because P(A) + P(B) = 1 where A is the event of getting X to be 0 or 1; while B is the event of getting X > 1. One or the other must happen. The events are complementary. We can rearrange that previous equation into P(B) = 1 - P(A).